natalie spent $135 on new shorts and tops. the shorts were 15% off and the tops were 20% off. if she saved a total of $24, how much did she spend on tops?
Answer by oberobic(2304)

t = $ spent on tops
s = $ spent on shorts
t + s = $135
So, t = $135-s
.
Define the discount equation:
.2t + .15s = $24
.
Multiply both sides by 100 to get rid of decimals
20t + 15s = 2400
.
Substitute t = 135-s
.
20(135-s) + 15s = 2400
2700 - 20s + 15s = 2400
.
Collect and simplify
-5s = -300
s = 60
.
Substitute s=60 to find t.
.
t + s = 135
t + 60 = 135
t = 75
.
Check using the discount equation.
.
.2t + .15s = 24 ??
.2(75) = 15
.15(60) = 9
15+9 =24
Correct.
.
Answer:
How much did she spend on tops? $75.
.
Done